OpenStudy (anonymous):

Simplify and write the trigonometric expression in terms of sine and cosine: Cosusecu/tanu=f(u)/g(u)? f(u)= g(u)=

OpenStudy (dumbcow):

sec = 1/cos so cos*sec = 1

geerky42 (geerky42):

Ok, like @dumbcow said, \(\cos\theta\sec\theta=1\), so you now have \(\dfrac{1}{\tan\theta}\), right? Just remember that \(\tan\theta = \dfrac{\sin\theta}{\cos\theta}\).

OpenStudy (anonymous):

ok i see. so what would i put for f(u) and g(u)? when i plugged in 1 for f(u) it was wrong and when i plugged in sinu/cosu for g(u) it also was wrong. Did i do this wrong?

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

It's because you still need to simplify further. You have \(\dfrac{1}{\tan\theta} = \dfrac{1}{\dfrac{\sin\theta}{\cos\theta}}\) Know what to do next?

OpenStudy (anonymous):

yeah you get costheta/ sintheta right?

OpenStudy (anonymous):

okay yeah this was correct thank you so much!

OpenStudy (dumbcow):

lol i cant read